In a sample of 500 persons from Town A, 200 are found to be consumers of Wheat. In a sample of 400 from Town B, 220 are found to be consumers of Wheat. Do these data reveal a significant difference between Town A and Town B as far as the proportion of Wheat consumers is concerned?


1
Expert's answer
2021-01-05T15:00:48-0500

H0: Two towns do not differ much as far as the proportion of wheat consumption. P1 = P2

H1: P1 ≠ P2

The Statistic test is

"Z = \frac{p_1-p_2}{\sqrt{PQ(\frac{1}{n_1}+ \frac{1}{n_2})}} \\

p_1 = \frac{200}{500} =0.4 \\

p_2 = \frac{220}{400} = 0.55 \\

P = \frac{n_1p_1+n_2p_2}{n_1+n_2} \\

= \frac{500(0.4) + 400(0.55)}{500+400} \\

= \frac{200+220}{900} \\

= 0.466 \\

Q = 1 -P = 1 -0.466 = 0.534 \\

Z = \frac{0.4-0.55}{\sqrt{(0.466)(0.534)(\frac{1}{500}+ \frac{1}{400})}} \\

= \frac{0.15}{0.03346} \\

= 4.48"

Calculated value Z = 4.48

Tabulated value of Z at 5% level of significance for a two tail test is 1.96

Calculated value >Tabulated value, H0 is rejected.

Hence the data reveal a significant difference between town A and town B so far as the proportion of wheat consumers is concerned.